OpenStudy (anonymous):

The product of 9 and a number is 12 less than three times that number. What is the number?

OpenStudy (anonymous):

product is the result of multiplying and " is " means equals 9n = 3n - 12 -- subtract 3n from both sides 9n - 3n = -12 6n = -12 -- divide by 6 n = -12/6 n = -2

OpenStudy (anonymous):

thank you so much you recive a meadl thanks

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

so its -2 or 2

OpenStudy (anonymous):

its -2 (negative 2)